OpenStudy (anonymous):

A two digit number is such that the product of its digits is 35. If 18 is added to the number, the digits interchange the places. Find the number. Solve using system of equations in two variables.

OpenStudy (dumbcow):

easy :P

OpenStudy (anonymous):

how

OpenStudy (dumbcow):

@FoolForMath will show you

OpenStudy (anonymous):

Lets say the number is \( 10x+y\) \[x\times y = 35 \tag{1} \] \[ 10x+y +18 = 10y+x \implies 18 + 9 x - 9 y =0 \tag{2} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I too got the same equation, but you will end up with a quadratic equation. I tried solving it, but I got a fractional answer.

OpenStudy (anonymous):

y=7 and x =5 .. the number is 57

OpenStudy (anonymous):

Can you show me how you got it?

OpenStudy (callisto):

ab = 35 10a+b+18 = 10b+a 10a +(35/a) +18 = 10(35/a) +a 9a^2 + 18a -315 =0 a=5 or a=-7 (rejected) b=35/5 = 7 => number is 57

OpenStudy (anonymous):

You got a = 5 by solving the quadratic equation?

OpenStudy (anonymous):

I must have made a mistake somewhere, thanks anyway.

OpenStudy (anonymous):

Who do I give the medal to? Fool or Callisto?

OpenStudy (callisto):

Fool :)

OpenStudy (anonymous):

Callisto!

OpenStudy (anonymous):

OK, Fool can give to Callisto, I'll give to Fool.

OpenStudy (dumbcow):

ok here is easy way...no need for equations what are the factors of 35.....5 and 7 so its either 57 or 75 75-57 = 18 answer = 57

OpenStudy (callisto):

using system of equations :P

OpenStudy (anonymous):

@dumbcow: There is another easy way, just go through the options ;)

OpenStudy (dumbcow):

well yeah if its multiple choice...but thats no fun

OpenStudy (callisto):

The easiest way is that the answer pops out immediately :)

OpenStudy (anonymous):

You're right callisto but this is a subjective question.